verify that each given function is a solution of the

differential equation.

8.y′′′′\:+\:4y′′′\:+\:3y\:=\:t;\:y\left(t\right)\:=\frac{t}{3}


1
Expert's answer
2021-09-03T07:34:06-0400
"y''''+4y'''+3y=t, y_1(t)=\dfrac{t}{3}"

"y=\dfrac{t}{3}=>y'=\dfrac{1}{3}, y''=y'''=y''''=0"

Substitute


"0+4(0)+3(\dfrac{t}{3})=t"




"t=t, True\ for\ t\in(-\infin, \infin)"

The function "y_1(t)=\dfrac{t}{3}" is a solution of the given differential equation.




"y''''+4y'''+3y=t, y_2(t)=e^{-t}+\dfrac{t}{3}"




"y=e^{-t}+\dfrac{t}{3}=>y'=-e^{-t}+\dfrac{1}{3}, y''=e^{-t},"




"y'''=-e^{-t},y''''=e^{-t}"

Substitute


"e^{-t}+4(-e^{-t})+3(e^{-t}+\dfrac{t}{3})=t"




"t=t, True\ for\ t\in(-\infin, \infin)"


The function "y_2(t)=e^{-t}+\dfrac{t}{3}" is a solution of the given differential equation.